Please help me solve these i need them today
145-5x=4x+10
2x+11=-16-7x
-15x-7x=10x+19
14-8x=34-15x
219-2x=34+19

Answer by Alan3354(69443)

145-5x=4x+10
Get all x terms on one side.
145-5x=4x+10
Add 5x
145 = 9x + 10
Get everything else on the other side.
Subtract 10
135 = 9x
Divide by 9
15 = x
x = 15
=====================
These are the same as the 1st one.
----
2x+11=-16-7x
-15x-7x=10x+19
14-8x=34-15x
219-2x=34+19